# Flexible Process Variant Binding in Information Systems with Software Product Line Engineering

### v1

@@ 2.7 FeatureIDE
FeatureIDE [51] is an Eclipse-based integrated development environment (IDE) for SPLs. It covers the four phases of Software Product Line Engineering set out in Section 2.6). To support domain analysis, FeatureIDE provides a graphical editor to model features and their dependencies in a feature model. Feature models contain constraints, are stored as XML files, and can be imported to or exported from FeatureIDE. Furthermore, FeatureIDE provides the user with convenient refactoring tools and inconsistency detection.
@@ 3 Previous Work
In our previous work, we introduced the approach of Process-Aware Information System Product Line (PAIS Product Line) [24]. This approach applies the concepts of SPL Engineering to the development and maintenance of similar PAISs in order to reduce development efforts and costs by enabling the reuse of common core artefacts.
@@ 3.1 Variation Points
A PAIS Product Line constitutes a common set of core artefacts including a core process model from which PAIS products may be derived. Various activities of the core process model may be declared as variation points, i.e. different implementations may be selected for these activities during PAIS Product derivation. In BPMN 2.0444https://www.omg.org/spec/BPMN/2.0/PDF, an activity corresponds to a step on which work is performed, which is either atomic or compound. While call activities (call subprocess, which itself comprise activities) constitute compound activities, tasks are atomic activities that cannot be further broken down. For the sake of this approach, the BPMN 2.0 send task, script task, service task, and business rule task are regarded as automated tasks as they all represent an automatic step in a business process without user interaction.
@@
Figure 2: Example Business Process [24]
@@ 3.2 Variability Modelling
Figure 3: Process Feature Model [24]
@@ 3.3 Feature Interaction
Each activity of a business process has input and output data whose data structure is specified at design time. The term process data structure is used to refer to the combined input and output data of all activities of a business process. Thus, the input/output data structure of an activity determines that part of the process data structure to which the activity has read/write access. Two activities with the same output data structure have the same write access. Hence, an activity that has the same output data structure as its preceding activity might overwrite the data of its predecessor, i.e. the data of the predecessor is lost. During requirements engineering, the order of the activities as well as their output data structure (i.e. write access) is specified. Therefore, the behaviour of the activities can be foreseen and data will not inadvertently be overwritten. Aside from passing on and accepting data, there is no interaction between the activities as they are executed in a predefined order and not in parallel. Consequently, there are no unintended feature interactions in PAIS Product Lines.
@@ 4.1 Requirements
In [24], we used the business process special parking permit for craftspersons in a simplified version as an example, which we deduced from a cooperation with German municipalities. In this work we extend the example towards a more complete version of the business process. Example 4.1 describes the business process.
@@
While the described control flow is common across all municipalities, the implementation of the business process varies. In some municipalities, the application check is carried out by a municipal clerk whereas in other municipalities the application is checked automatically by comparing the data of the application with the data of the craftsperson who is officially registered with the authorities. Besides, when the application is rejected the applicant may be notified via e-mail, or sms, or by a municipal clerk depending on the municipality and the choice of the applicant. The applicant might choose the way she will be notified from a list of available notification means.
Furthermore, the process data structure also varies as the municipalities require the applicant to fill in different data. In some municipalities, the parking permit is valid for every car whereas in others the applicant needs to provide the number plate for which the parking permit shall be valid in order to prevent misuse of the parking permit for the craftsperson’s private car. In addition, to be able to automatically check the application, the craftsperson needs to provide the commercial register number under which she is registered with the authorities.
@@
In a nutshell, PAIS Product Lines need to comprise common core artefacts including a process model from which similar PAIS Products can be derived in order to reduce the development efforts in comparison to developing the similar PAIS Products separately. One or more implementations shall be selectable for an activity that constitutes a variation point. The implementations shall be selectable at compile time, start time, and runtime and the process data structure shall be customisable for each PAIS Product as well. The business process of checking an application for the special parking permit and the resulting requirements were studied in a cooperation with German municipalities. While in previous work we used a simplified version of the business process as an example, in this work we use a more comprehensive version and deduce the requirements for our enhanced approach accordingly. The enhanced approach of PAIS Product Lines shall meet the following requirements:
@@ 1. R1:
    PAIS Products need to be derivable from a PAIS Product Line which denotes a set of common core artefacts including a core process model in order to avoid the development of redundant software artefacts for similar PAISs.
@@ 2. R2:
    Multiple implementations for one activity shall be selectable.
@@ 3. R3:
    During PAIS Product derivation the process data structure shall be adaptable allowing different organisation to use different process data structure in their derived PAIS Product.
@@ 6. R6:
    To enable selecting activity implementations during start and runtime based on input of users dynamic feature binding becomes necessary. When using dynamic feature binding, the challenges described in Section 2.5 need to be tackled:
@@ 4.2 Assumptions
For the approach to work, the following assumptions hold:
  • The control flow of the core process model included in a PAIS Product Line is specified and immutable, i.e. the process model does not change during PAIS Product derivation. Hence, all derived PAIS Products will have the same process model and thus an identical control flow. While this work focuses on implementations variability, other approaches deal with control flow variability [52, 21, 23, 22].
  • The concepts will be applied to activity-centric business processes. Object- and data-centric approaches to business process management like those discussed in [30] and [49] are excluded in this approach.
@@ 4.3 Multiple Implementations of an Activity
When allowing for multiple implementations of an activity, no execution order of these implementations is specified. If, for example, in Figure 5, for activity Check application both implementations Automatic Check and Manual Check are selected, both implementations run in parallel when the process instance passes this activity. As the implementations of an activity may be executed in isolation there is no unintended technical feature interaction. However, implementations of one activity may influence each other via the data they write. Due to non-determinism it might not been foreseeable which implementation finishes last and thereby overwrites the data of its predecessors, i.e. the results are unpredictable. Consider Example 4.3.
@@
(Feature Interactions in PAIS Product Lines): In Figure 5, activity Check application writes the decision on whether an application is justified to the process data structure. Consequently, both implementations may write the decision as well, i.e. the application is checked both automatically and by a municipal clerk. When selecting both implementations (i.e. manual and automatic) for activity Check application, the implementation that finishes last determines whether the application is justified. For example, the result of the automatic application check is overwritten by the decision of the municipal clerk if he/she takes longer than the automatic application check. This leads to unpredictable results depending on which implementation writes the data last. This can be seen as unintended feature interaction.
@@
Additional code can be included when two implementations write to the same data structure. For example, additional code could determine that the application is not justified as soon as at least one of the two implementations decides that the application is not justified. This additional code aggregates the data from the different implementations. We refer to this additional code as aggregation code. Another aggregation code might constitute a majority vote. If there are multiple application check implementations, the application is rejected when a majority of the implementations consider the application being unjustified. As opposed to derivative modules, aggregation code needs to be placed in the process feature model as aggregation code does not solely resolve technical dependencies. Aggregation code introduces resolution on a business level. There might be multiple aggregation codes from which one may be selected during PAIS product derivation. This choice needs to be indicated in the process feature model. In the process feature model all implementations of one activity are grouped as well as all available aggregation codes for that activity. A dashed arrow introduces the constraint that one aggregation code needs to be selected when more than one implementation of an activity is selected. Consider the following example.
@@
which reads as follows: If the number of selected implementations is greater than one, one aggregation code implementation needs to be selected, i.e. if two items from Group I are selected, one item from Group A needs to be co-selected as well. The dashed arrow, which represents a conditional requires, was taken from literature (cf. Section 2.1) and supplemented with a label.
@@ 4.4 Varying Data Structures
In the following, the process of checking an application for the special parking permit (cf. Examples 4.1 and 4.4) is used to explain the concept of varying data structures. Figure 6 shows the process data structure associated with Example 4.4 as UML class diagram. The application corresponds to a composition of the data of the applicant, the data of the company, and the car information, whereas the latter and the field commercial register number in the data of the company are optional (indicated with question marks). The optional parts of the data structures are only included if necessary.
@@ 4.5 Static and Dynamic Feature Binding
The activity implementations are independent from each other and have no technical inter-dependencies. The single point of interaction among the activity implementations is the write and read access to the process data structure. Consequently, the order in which the activity implementations are bound does not influence their behaviour.
If an activity implementation runs during process execution, as it was bound during process start and selected by a user, the implementation will not be unbound during execution. Once selected for a process instance, an activity implementation will always finish it execution. Hence, no data will be lost.
Shared data and system processes do not pose a problem as activity implementations are independent and have no technical dependencies. Furthermore, activity implementations may not be unbound during process execution once they are selected.
5 Proof of Concept
To validate the presented approach, Example 4.1 is implemented as a proof of concept. The Camunda Platform is used as Workflow Management System. [^5] https://docs.camunda.org/manual/7.19/ First, the properties are stated that need to be provided by the proof of concepts. Then, from a technical perspective the proof of concept shall show how to customize the process data structure during PAIS Product derivation, how to bind multiple implementations for one activity, and how to bind the implementations at different times. Finally, it is portrayed how the development and maintenance may be supported by a tool.

### v2

@@ 2.7 FeatureIDE
FeatureIDE (Thüm et al., 2014) is an Eclipse-based integrated development environment (IDE) for SPLs. It covers the four phases of Software Product Line Engineering set out in Section 2.1. To support domain analysis, FeatureIDE provides a graphical editor to model features and their dependencies in a feature model. Feature models contain constraints, are stored as XML files, and can be imported to or exported from FeatureIDE. Furthermore, FeatureIDE provides the user with convenient refactoring tools and inconsistency detection.
@@ 3 Previous Work
In our previous work, we introduced the approach of Process-Aware Information System Product Line (PAIS Product Line) (Hehnle and Reichert, 2023). This approach applies the concepts of SPL Engineering to the development and maintenance of similar PAISs in order to reduce development efforts and costs by enabling the reuse of common core artefacts.
@@ 3.1 Variation Points
A PAIS Product Line constitutes a common set of core artefacts including a core process model from which PAIS products may be derived. Various activities of the core process model may be declared as variation points, i.e. different implementations may be selected for these activities during PAIS Product derivation. Throughout this work, BPMN 2.0555https://www.omg.org/spec/BPMN/2.0.2/PDF is used as notation to model business processes as we agree with other researchers that BPMN is the de-facto industry standard for process modelling (Döhring and Zimmermann, 2011; Hongyan Zhang et al., 2014; Delgado et al., 2022). In addition, BPMN process models can be executed by a process engine, which is harnessed by our proof-of-concept. In BPMN 2.0, an activity corresponds to a step on which work is performed, which is either atomic or compound. While call activities (call subprocess, which itself comprise activities) constitute compound activities, tasks are atomic activities that cannot be further broken down. For the sake of this approach, the BPMN 2.0 send task, script task, service task, and business rule task are regarded as automated tasks as they all represent an automatic step in a business process without user interaction.
@@
Figure 2: Example Business Process (Hehnle and Reichert, 2023)
@@ 3.2 Variability Modelling
Figure 3: Process Feature Model (Hehnle and Reichert, 2023)
@@ 3.3 Feature Interaction
Each activity of a business process has input and output data whose data structure is specified at design time. The term process data structure is used to refer to the combined input and output data of all activities of a business process. Thus, the input/output data structure of an activity determines that part of the process data structure to which the activity has read/write access. Two activities with the same output data structure have the same write access. Hence, an activity that has the same output data structure as its preceding activity might overwrite the data of its predecessor, i.e. the data of the predecessor is lost. During requirements engineering, the order of the activities as well as their output data structure (i.e. write access) is specified. Therefore, the behaviour of the activities can be foreseen and data will not inadvertently be overwritten. Aside from passing on and accepting data, there is no unintended interaction between the activities as they are executed in a predefined order. If two activities in a process are modelled in parallel (e.g. after a gateway), they are often still executed sequentially by common Workflow Management Systems (WfMS) by default. When configured accordingly, parallel modelled activities may be executed in different threads to achieve concurrency. While there are some ways in which the threads may interact from a technical perspective, e.g. accessing shared memory, pausing one thread in order to wait for the completion of another, these interactions would have to be developed explicitly and would therefore be intended. But, in general, these kinds of interactions do not reflect the requirements of a business process. Reading shared memory or other interactions between activities modelled in parallel (i.e. executed in parallel threads) predominantly contradict the requirements of business processes. Activities of a business process correspond to logically isolated features, which in general read process data, transform the data (e.g. by calculations or by prompting for user input) or call web services, and write the results back to the process data. Write access to process data needs to be considered during design time as mentioned previously. Consequently, there are no unintended feature interactions in PAIS Product Lines.
@@ 4.1 Requirements
The requirements for our approach were deduced from a cooperation with German municipalities. German municipalities have organised various events at regional and state level (See Table 4.1 for events we attended) to share experiences in the digitisation of municipal business processes involving various stakeholders, including the respective state government, digitisation companies from the public and private sectors, and municipal umbrella organisations. During these events, experiences are exchanged on approaches and technical, legal, and financial challenges in the digitisation of municipal business processes. Due to our knowledge in business process management and digitisation, one co-author and two co-workers [^6] The referenced co-author simultaneously held and holds part-time positions as a research associate and as an employee at a digitisation company. The mentioned co-workers were employed by the same firm. were invited to attend three events (cf. Table 4.1) and act as facilitators during the breakout sessions. In the following, we describe the three events and our role.
In Event 1, there was a one breakout session workshop (90 minutes), in which a task force of three municipalities discussed the requirements of the business process special parking permit for craftspersons. The objective of the task force was to agree on common requirements for three business process so that uniform PAISs could be developed while reducing redundancies and costs. Each municipality was supposed to digitise one business process and share it with the partner municipalities, so that each municipality only had to digitise one business process but could use three. We acted as facilitators in the breakout session in which the requirements for the business process special parking permit for craftspersons should be aligned.
A variety of collaboration and participation approaches have been proposed to elicit requirements and to model business processes in joint sessions by involving different stakeholders (e.g. end users) to include diverse perspectives (Barjis, 2011; Lai et al., 2014; Front et al., 2017). While these approaches focus on different methods with varying level of detail (e.g. incorporating group storytelling or role plays), all have joint sessions in common in which domain experts present the requirements verbally, a facilitator mediates between them, and a modelling expert designs the business processes. Although we did not apply any specific approach, we followed the basic idea of these approaches. As opposed to the mentioned approaches, we were facing the challenge to not only involve different stakeholders of the same business process but rather stakeholders of multiple similar yet slightly different business processes.
In line with the roles described by Lai et al. (2014), at least one domain expert (i.e. municipal clerk who is a process participant) and an IT expert from each municipality were present to share the expectations and activities of the business process. We acted as facilitators and process modellers to collect the requirements and model the business process graphically. Our main responsibility was to mediate between the different requirements of the municipalities, reach an agreement and thereby model a standardised business process. While available software tools (Barjis, 2011; Front et al., 2017) or even interactive tabletops (Döweling et al., 2013) support collaborative and participative process modelling, we relied on whiteboards and sticky notes like Lai et al. (2014). While we acknowledge that under normal circumstances 90 minutes are not sufficient to elicit the requirements for a business process, in this case, the three municipalities already had elicited the requirements of their respective business process. Consequently, during the session, the requirements of the three municipalities only had to be collected, compared, and synthesised, and areas of alignment and divergence had to be identified. After the event, we tried to make compromise proposals by phone and email. However, despite our efforts, all compromise proposals were rejected without further workshops. Consequently, the differences in the requirements proved impossible to develop a common PAIS.
Significantly more municipalities attended Event 2 to align their digitisation strategies and form collaborations at the state level. While the event had a fixed agenda with presentations (e.g. by a representative of the interior ministry of the corresponding state government), the event was designed to provide room for networking and informal exchanges between municipalities, government institutions, and private companies supporting the digitisation endeavours of the municipalities and state government. During informal conversations with further municipalities, we could confirm that the requirements regarding implementation variability of business processes we deduced for the special parking permit also apply to other municipal business processes.
Finally, in breakout sessions during Event 3, once again we acted as mediators and process designers to collect and compare the requirements of other municipal business processes, which include enrolling children in kindergarten and applying and issuing the civil status documents for a citizen.
In summary, we could specifically elicit the requirements for the business process special parking permit for craftspersons of three German municipalities and deduce the challenge of implementation variability of business processes. Furthermore, we could confirm that this challenge is not restricted to this business process but can be applied to other municipal business processes.
ID D1 Important Agenda Topics Participants
1 8 Exchange of digitisation approaches Legal challenges Technical challenges Discussions about joint digitisation of concrete business processes in break-out sessions in smaller groups2 17 municipal employees (8 municipalities) 2 publicly owned corporations for digitisation 2 representatives of the interior ministry of the state government 1 co-author and 1 co-worker as mediator 3 representatives of another digitisation company
2 6 Legal and financial conditions Discussion about future collaboration between municipalities, companies, and state government 3 breaks for informal exchanges around 150 workshop participants including representatives from municipalities, companies (including co-author and co-worker as guests), publicly owned corporations for digitisation, and state government
3 6 Benefits of digitisation Prioritisation for the digitisation of municipal business processes Break-out sessions to discuss commonalties and differences of concrete municipal business processes2 14 municipal employees (10 municipalities) 1 employee of a municipal umbrella organisation 1 employee of a publicly owned corporation for digitisation 1 co-author and 2 co-workers as mediators and process designers
  • 1
    Duration in hours excluding preparing and cleaning up the meeting rooms but including the coffee and lunch break, in which the topics were informally discussed.
  • 2
    Collaborative requirements elicitation facilitated by co-author/co-workers
Events for Digitising Municipal Business Processes
Example 4.1 describes the business process special parking permit for craftspersons according to the requirements elicited during the conducted workshop, which, according to our findings, represent general widespread challenges in the digitisation of business processes in German municipalities. In (Hehnle and Reichert, 2023), we used the business process in a simplified version as an example.
@@
While the described control flow is common across all municipalities, the implementation of the business process varies. In some municipalities, the application check is carried out by a municipal clerk whereas in other municipalities the application is checked automatically by comparing the data of the application with the data of the craftsperson who is officially registered with the authorities. Besides, when the application is rejected the applicant may be notified via e-mail, or SMS, or by a municipal clerk depending on the municipality and the choice of the applicant. The applicants might choose the way they will be notified from a list of available notification means.
Furthermore, the process data structure also varies as the municipalities require the applicant to fill in different data. In some municipalities, the parking permit is valid for every car whereas in others the applicant needs to provide the number plate for which the parking permit shall be valid in order to prevent misuse of the parking permit for the craftsperson’s private car. In addition, to be able to automatically check the applications, the craftspersons need to provide the commercial register numbers under which they are registered with the authorities.
@@
In a nutshell, PAIS Product Lines need to comprise common core artefacts including a process model from which similar PAIS Products can be derived in order to reduce the development efforts in comparison to developing the similar PAIS Products separately. One or more implementations shall be selectable for an activity that constitutes a variation point. The implementations shall be selectable at compile time, start time, and runtime and the process data structure shall be customisable for each PAIS Product as well. The enhanced approach of PAIS Product Lines shall meet the following requirements:
@@ 1. R1:
    In line with our previous work, PAIS Products need to be derivable from a PAIS Product Line which denotes a set of common core artefacts including a core process model in order to avoid the development of redundant software artefacts for similar PAISs.
@@ 2. R2:
    Multiple implementations for one activity shall be selectable. Only the selected implementations shall be contained in the derived PAIS to keep the size of the latter to a minimum.
@@ 3. R3:
    During PAIS Product derivation the process data structure shall be customisable allowing different organisation to use different process data structure in their derived PAIS Product.
@@ 6. R6:
    To enable selecting activity implementations during start and runtime based on input of users’ dynamic feature binding becomes necessary. When using dynamic feature binding, the challenges described in Section 2.3 need to be tackled:
@@ 4.2 Assumptions
Compared to previous work (Hehnle and Reichert, 2023), we have lifted all but two assumptions:
  • The control flow of the core process model included in a PAIS Product Line is specified and immutable, i.e. the process model does not change during PAIS Product derivation. Hence, all derived PAIS Products will have the same process model and thus an identical control flow. While this work focuses on implementations variability, other approaches deal with control flow variability (van der Aalst et al., 2006; Hallerbach et al., 2008, 2010b, 2010a). The requirements we elicited with the three German municipalities for the business process special parking permit do not require control flow variability. Nevertheless, based on informal discussions with further German municipalities we are aware of cases in which both implementation variability and control flow variability of business processes are necessary. However, we expect to solve lots of cases using implementation variability, only. Future work shall reveal how to combine implementation variability and control flow variability to cover all cases of business process variability.
  • The concepts will be applied to activity-centric business processes. There are different approaches to business process management with each having advantages and disadvantages. While object- and data-centric approaches like those discussed by Künzle and Reichert (2009) and Steinau et al. (2019) have some advantages, we will apply our concepts to BPMN as the de-facto industry standard (i.e. activity-centric business processes) to achieve broad acceptance.
@@ 4.3 Multiple Implementations of an Activity
Building on previous work, PAIS Products are derived from a PAIS Product Line consisting of a set of common core artefacts including the process model and the activity implementations (cf. Requirement R1). When allowing for multiple implementations of an activity, no execution order of these implementations is specified. If, for example, in Figure 5, for activity Check application both implementations Automatic Check and Manual Check are selected, both implementations run in parallel when the process instance passes this activity. As the implementations of an activity may be executed in isolation there is no unintended technical feature interaction. However, implementations of one activity may influence each other via the data they write. Due to non-determinism, it might not be foreseeable which implementation finishes last and thereby overwrites the data of its predecessors, i.e. the results are unpredictable. Consider Example 4.3.
@@
(Feature Interactions in PAIS Product Lines): In Figure 5, activity Check application writes the decision on whether an application is justified to the process data structure. Consequently, both implementations may write the decision as well, i.e. the application is checked both automatically and by a municipal clerk. When selecting both implementations (i.e. manual and automatic) for activity Check application, the implementation that finishes last determines whether the application is justified. For example, the result of the automatic application check is overwritten by the decision of the municipal clerk if the clerk takes longer than the automatic application check. This leads to unpredictable results depending on which implementation writes the data last. This can be seen as unintended feature interaction.
@@
Additional code can be included when two implementations write to the same data structure. For example, additional code could determine that the application is not justified as soon as at least one of the two implementations decides that the application is not justified. This additional code aggregates the data from the different implementations. We refer to this additional code as aggregation code. Another aggregation code might constitute a majority vote. If there are multiple application check implementations, the application is rejected when a majority of the implementations consider the application being unjustified. As opposed to derivative modules, aggregation code needs to be placed in the process feature model as aggregation code does not solely resolve technical dependencies. Aggregation code introduces resolution on a business level. There might be multiple aggregation codes from which one may be selected during PAIS product derivation. This choice needs to be indicated in the process feature model. In the process feature model, all implementations of one activity are grouped as well as all available aggregation codes for that activity. A dashed arrow introduces the constraint that one aggregation code needs to be selected when more than one implementation of an activity is selected. Consider the following example.
@@
which reads as follows: If the number of selected implementations is greater than one, one aggregation code implementation needs to be selected, i.e. if two items from Group I are selected, one item from Group A needs to be co-selected as well. The dashed arrow, which represents a conditional requires, was taken from literature (cf. Section 2.2) and supplemented with a label. Using aggregation code, it becomes possible to prevent unintended interaction when selecting multiple implementations for one activity (cf. RequirementsR2 and R5).
@@ 4.4 Varying Data Structures
In the following, the process of checking an application for the special parking permit (cf. Examples 4.1 and 4.4) is used to explain the concept of varying data structures. Figure 6 shows the process data structure associated with Example 4.4 as UML class diagram. The application corresponds to a composition of the data of the applicant, the data of the company, and the car information, whereas the latter and the field commercial register number in the data of the company are optional (indicated with question marks). The optional parts of the data structures are only included if necessary (cf. Requirement R3).
@@
Using the feature model presented in Figure 7, the constraints between the implementations are the optional data structure fields become apparent (cf. Requirement R4).
@@ 4.5 Static and Dynamic Feature Binding
The activity implementations are independent from each other and have no technical inter-dependencies. The single point of interaction among the activity implementations is the write and read access to the process data structure. Consequently, the order in which the activity implementations are bound does not influence their behaviour (cf. Requirement R6c).
If an activity implementation runs during process execution, as it was bound during process start and selected by a user, the implementation will not be unbound during execution. Once selected for a process instance, an activity implementation will always finish it execution. Hence, no data will be lost (cf. Requirement R6a).
Shared data and system processes do not pose a problem as activity implementations are independent and have no technical dependencies. Furthermore, activity implementations may not be unbound during process execution once they are selected (cf. Requirement R6b).
5 Proof-of-Concept
To validate the presented approach, Example 4.1 is implemented as a proof-of-concept, which is available at GitHub [^7] https://github.com/hehnle/process-implementation-variability. For the proof-of-concept, an unaltered version of Camunda Platform [^8] https://docs.camunda.org/manual/7.22/ is used as Workflow Management System (WfMS) embedded in a Java Spring Boot application. In an effort to avoid proprietary software for our approach and guarantee seamless updates of Camunda Platform we did not customise Camunda Platform or other tools used for the proof-of-concept. When customising tools directly or indirectly by developing plugins for a tool, these customisations and plugins need to be kept up to date and compatible with said tool. As we are not able to maintain software sustainably, we solely use available tools as they are.
In the following, the properties are stated that need to be provided by the proof-of-concept. Then, from a technical perspective the proof-of-concept shall show how to customize the process data structure during PAIS Product derivation, how to bind multiple implementations for one activity, and how to bind the implementations at different times. Finally, it is portrayed how the development and maintenance may be supported by a tool.